OpenStudy (anonymous):

The temperature of a chemical reaction ranges between −10°C and 50°C. The temperature is at its lowest point when t = 0 and completes one cycle over a six-hour period. What is a sine function that would model this reaction? f(x) = 30 sin pi over 3x + 20 f(x) = 60 sin pi over 3x + 30 f(x) = 30 sin 6x + 20 f(x) = 60 sin 6x + 30

OpenStudy (anonymous):

w=2pi/T->w=2pi/6=pi/3 then(A), since 30+20=50 and -30+20=-10